# Class-RAG: Content Moderation with Retrieval Augmented Generation

### v1

@@ 5.4 Instruction Following Ability
The instruction following ability of a large language model (LLM) refers to its capacity to comprehend and accurately respond to given instructions. In this section, we investigate the ability of Class-RAG to follow the guidance from reference examples and generate responses consistent with these examples. To evaluate this, we utilized the ID test set with a flipped ID library, which contains the same examples as the original ID library but with flipped labels ("unsafe" → "safe", "safe" → "unsafe") and removed explanations. The results, presented in Table 5, demonstrate that Class-RAG possesses a strong instruction following ability. Notably, the predicted labels of 99.49% of ground-truth safe examples were successfully flipped from "safe" to "unsafe", while the predicted labels of 12.29% of ground-truth unsafe examples were flipped from "unsafe" to "safe". This disparity in flipping ratios between ground-truth safe and unsafe examples can be attributed to the safety fine-tuning of the Llama3 model, which has been designed to prevent generating harmful responses and has memorized unsafe content.
@@
5.5 Performance Scalability with Retrieval Library Size
We conducted an investigation to examine the impact of retrieval library size on the performance of Class-RAG, with results presented in Table 4. Specifically, we utilized the in-distribution (ID) library collected from the CoPro training set and the external (EX) library collected from the validation sets of I2P++ and UD. To assess the effect of library size on performance, we created downscaled versions of the external library, denoted as EX ($\frac{1}{8}$), EX ($\frac{1}{4}$), and EX ($\frac{1}{2}$), which were constructed by re-clustering the full external library to $\frac{1}{8}$, $\frac{1}{4}$, and $\frac{1}{2}$ of its original size, respectively.
Our results demonstrate that model performance consistently improves with increasing retrieval library size. On the I2P++ dataset, we observed AUPRC scores of 0.235, 0.501, 0.554, 0.577, and 0.677 when adding 0, $\frac{1}{8}$, $\frac{1}{4}$, $\frac{1}{2}$, and the full size of the external library, respectively. Similarly, on the UD dataset, AUPRC scores increased from 0.914 to 0.959, 0.963, 0.964, and 0.974 with the addition of 0, $\frac{1}{8}$, $\frac{1}{4}$, $\frac{1}{2}$, and the full size of the external library, respectively.
@@ 5.5 Performance Scalability with Retrieval Library Size
5.6 Performance Scalability with Reference Example Number
@@
We conducted a further investigation to examine the impact of the number of reference examples on the performance of Class-RAG. Specifically, we evaluated the model’s performance when adding 0, 2, 4, 6, and 8 reference examples, with an equal number of safe and unsafe examples added in each case. The results, presented in Table 6, demonstrate that the performance of Class-RAG consistently improves with the addition of more reference examples. On the I2P++ dataset, we observed average AUPRC scores of 0.303, 0.632, 0.677, 0.715, and 0.721 when using 0, 2, 4, 6, and 8 reference examples, respectively. Similarly, on the UD dataset, average AUPRC scores increased from 0.932 to 0.965, 0.974, 0.978, and 0.980 with the addition of 0, 2, 4, 6, and 8 reference examples, respectively.
@@
While our results indicate that performance improves with the number of reference examples, we also observe that this improvement becomes saturated at around 8 reference examples. Furthermore, adding more reference examples incurs a higher computational cost compared to scaling up the retrieval library size. Therefore, while increasing the number of reference examples can enhance performance, it is essential to balance this with the associated computational expense.
Table 6: AUPRC scores for Class-RAG on the I2P++ and UD external datasets using different numbers of reference examples. Higher AUPRC scores indicate better performance
Obfuscations 0 ref. 2 ref. 4 ref. 6 ref. 8 ref.
I2P++
None 0.377 0.795 0.791 0.838 0.839
change_case 0.360 0.824 0.843 0.873 0.870
insert_punctuation_chars 0.227 0.292 0.318 0.332 0.354
insert_text 0.369 0.810 0.816 0.856 0.854
whitespace_chars 0.284 0.515 0.601 0.648 0.673
merge_words 0.368 0.807 0.815 0.859 0.856
replace_similar_chars 0.202 0.422 0.549 0.540 0.540
simulate_typos 0.236 0.708 0.742 0.788 0.779
split_words 0.305 0.516 0.613 0.701 0.724
AVERAGE 0.303 0.632 0.677 0.715 0.721
UD
None 0.959 0.984 0.985 0.991 0.991
change_case 0.956 0.988 0.991 0.994 0.993
insert_punctuation_chars 0.900 0.911 0.931 0.934 0.943
insert_text 0.951 0.987 0.988 0.992 0.992
whitespace_chars 0.933 0.953 0.971 0.976 0.979
merge_words 0.952 0.989 0.990 0.994 0.993
replace_similar_chars 0.896 0.934 0.953 0.959 0.960
simulate_typos 0.917 0.979 0.984 0.988 0.987
split_words 0.928 0.961 0.972 0.980 0.984
AVERAGE 0.932 0.965 0.974 0.978 0.980
5.7 Performance with Different Embedding Models
@@
The choice of embedding model is crucial for retrieving relevant content in our proposed approach. In this section, we investigate the impact of two different embedding models on the performance of Class-RAG: DRAGON RoBERTa Lin et al. (2023) and WPIE. DRAGON is a bi-encoder dense retrieval model that embeds both queries and documents into dense vectors, enabling efficient search for relevant information from a large number of documents. We utilize the context encoder component of DRAGON in our experiments. In contrast, WPIE is a 4-layer XLM-R Conneau et al. (2020) model that has been pre-trained on safety data, producing two outputs: an unsafe probability and a prompt embedding.
@@
Our results, presented in Table 3, demonstrate that the DRAGON RoBERTa embedding outperforms WPIE. Specifically, DRAGON RoBERTa achieves an average AUPRC of 0.938 on the ID test set and 0.920 on the OOD test set, surpassing the performance of WPIE, which obtains an average AUPRC of 0.912 on the ID test set and 0.901 on the OOD test set. Future work will involve exploring the effectiveness of additional embedding models to further enhance the performance of Class-RAG.

### v2

@@ 5.4 Instruction Following Ability
The instruction following ability of a LLM refers to its capacity to comprehend and accurately respond to given instructions. In this section, we investigate the ability of Class-RAG to follow the guidance from reference examples and generate responses consistent with these examples. It is crucial for Class-RAG to adapt its behavior to updates in the retrieval library. To evaluate this, we utilized the ID test set with a flipped ID library, which contains the same examples as the original ID library but with flipped labels ("unsafe" → "safe", "safe" → "unsafe") and removed explanations. The results, presented in Table 5, demonstrate that Class-RAG possesses a strong instruction following ability. Notably, the predicted labels of 99.49% of ground-truth safe examples were successfully flipped from "safe" to "unsafe", while the predicted labels of 12.29% of ground-truth unsafe examples were flipped from "unsafe" to "safe". This disparity in flipping ratios between ground-truth safe and unsafe examples can be attributed to the safety fine-tuning of the Llama3 model, which has been designed to prevent generating harmful responses and has memorized unsafe content.
@@
Figure 2: Impact of external retrieval library size (top) and reference example number (bottom) on average AUPRC. Detailed results are presented in Tables 6 and Table 7, respectively.
5.5 Impact of Retrieval Library Size
We investigated the impact of external retrieval library size on model performance, with results presented in Figure 2 and Table 6. To do this, we constructed new retrieval libraries by augmenting the in-distribution (ID) library with external libraries of varying sizes. The external (EX) library was sourced from the validation sets of I2P++ and UD. We created downscaled versions of the external library, denoted as EX ($\frac{1}{8}$), EX ($\frac{1}{4}$), and EX ($\frac{1}{2}$), which were constructed by re-clustering the full external library (1691 examples) to 212, 425, and 850 examples, respectively.
Our results show that model performance consistently improves with increasing external retrieval library size. On the I2P++ dataset, AUPRC scores increased from 0.235 to 0.677 as the external library size grew from 0 to 1691 examples. Specifically, we observed AUPRC scores of 0.501, 0.554, and 0.577 for external library sizes of 212, 425, and 850 examples, respectively. A similar trend was observed on the UD dataset, where AUPRC scores increased from 0.914 to 0.974 as the external library size increased.
@@ 5.5 Impact of Retrieval Library Size
5.6 Impact of Reference Example Number
@@
We conducted a further investigation to examine the impact of the number of reference examples on the performance of Class-RAG. Specifically, we evaluated the model’s performance when adding 0, 2, 4, 6, and 8 reference examples, with an equal number of safe and unsafe examples added in each case. The results, presented in Figure 2 and Table 7, demonstrate that the performance of Class-RAG consistently improves with the addition of more reference examples. On the I2P++ dataset, we observed average AUPRC scores of 0.303, 0.632, 0.677, 0.715, and 0.721 when using 0, 2, 4, 6, and 8 reference examples, respectively. Similarly, on the UD dataset, average AUPRC scores increased from 0.932 to 0.965, 0.974, 0.978, and 0.980 with the addition of 0, 2, 4, 6, and 8 reference examples, respectively.
@@
While our results indicate that performance improves with the number of reference examples, we also observe that this improvement becomes saturated at around 8 reference examples. Furthermore, adding more reference examples leads to more input tokens and incurs a higher computational cost compared to scaling up the retrieval library size. Therefore, while increasing the number of reference examples can enhance performance, it is essential to balance this with the associated computational expense.
5.7 Impact of Embedding Models
@@
The choice of embedding model is crucial for retrieving relevant content in our proposed approach. In this section, we investigate the impact of two different embedding models on the performance of Class-RAG: DRAGON RoBERTa Lin et al. (2023) and WPIE (Whole Post Integrity Embedding) Meta (2021). DRAGON is a bi-encoder dense retrieval model that embeds both queries and documents into dense vectors, enabling efficient search for relevant information from a large number of documents. We utilize the context encoder component of DRAGON in our experiments. To investigate the impact of alternative embedding models on our approach, we also evaluate a variant of WPIE. The WPIE model we test is a 4-layer XLM-R Conneau et al. (2020) model that has been pre-trained on content moderation data, yielding two distinct outputs: an unsafe probability estimation and a prompt embedding representation.
@@
Our results, presented in Table 3, demonstrate that the DRAGON RoBERTa embedding outperforms WPIE. Specifically, DRAGON RoBERTa achieves an average AUPRC of 0.938 on the CoPro test set, surpassing the performance of WPIE, which obtains an average AUPRC of 0.912. Future work will involve exploring the effectiveness of additional embedding models to further enhance the performance of Class-RAG.